Should print DRAFT (in greek translation and greek letters) on a moved position.

\usepackage[timestamp,bottom,greek]{draftcopy}

Test 16 This is a test page 1

Test 16 This is a test page 2

This is a testpage 1 This is a testpage 2 This is a testpage 3 This is a testpage 4